OpenStudy (anonymous):

Simplify: (sin theta - cos theta)^2 + (sin theta + cos theta)^2

OpenStudy (anonymous):

\[(\sin \theta-\cos \theta)^2+(\sin \theta +\cos \theta)^2\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@joftheworld

OpenStudy (anonymous):

open upp squares and add up you will get it =2

OpenStudy (anonymous):

Thank you!!